# Parallel Program Analysis on Path Ranges

### v1

@@ Abstract
Symbolic execution is a software verification technique symbolically running programs and thereby checking for bugs. Ranged symbolic execution performs symbolic execution on program parts, so called path ranges, in parallel. Due to the parallelism, verification is accelerated and hence scales to larger programs.
In this paper, we discuss a generalization of ranged symbolic execution to arbitrary program analyses. More specifically, we present a verification approach that splits programs into path ranges and then runs arbitrary analyses on the ranges in parallel. Our approach in particular allows to run different analyses on different program parts. We have implemented this generalization on top of the tool CPAchecker and evaluated it on programs from the SV-COMP benchmark. Our evaluation shows that verification can benefit from the parallelisation of the verification task, but also needs a form of work stealing (between analyses) as to become efficient.
@@ 1 Introduction
In recent years, automatic software verification has become a mature technique reaching levels of industrial software. This progress is achieved by verification tools employing a bunch of different techniques for analysis, like predicate abstraction, bounded model checking, k-induction, property-directed reachability, or automata-based methods. As individual tools still have their specific strength and weaknesses, recent research in software verification has studied cooperative verfication [1]. In cooperative verification, tools cooperate on the task of verification. This principle has already been implemented in various forms [2, 3, 4, 5], in particular also as cooperations of testing and verification [6, 7, 8, 9].
@@
Cooperative verification often takes the form of a sequential or cyclic combination in which tools run in some sequential order, passing information from one to the next tool. Parallel combinations of different tools in the majority of cases employ portfolio approaches, simply running the different tools on the same task in parallel. The reason for this is often the lack of appropriate splitting techniques, which could divide the program into parts on which analyses can be run in parallel.
@@
The proposal of Siddiqui and Khurshid [10] for ranged symbolic execution provided exactly one such splitting technique. They defined so called path ranges as parts of the program on which several symbolic execution instances can be run in parallel. A path range describes a set of program paths defined by two inputs to the program, where the path triggered by the first input is the lower bound and the path for the second input is the upper bound of the range. All paths in between, i.e., paths such that (based on some ordering on paths), make up a range. Siddiqui and Khurshid experimentally confirmed that symbolic execution can benefit from this splitting into path ranges and this helps to scale the technique to larger programs.
@@
In this paper, we generalize ranged symbolic execution to arbitrary analyses. Our generalization concerns the use of analyses. Instead of its specialization to one analysis, namely symbolic execution, we present ranged analysis in which arbitrary analyses (in particular also different ones) can be run on the different ranges of a program in parallel. In this, we employ a form of work stealing in which one analysis can – once done with its own range – take over ranges assigned to other analyses as to better balance workloads. We have furthermore extended the approach of [10] by a new splitting technique. Instead of only choosing ranges randomly, we propose a splitting technique along loop bounds of arbitrary, user-configurable numbers . A loop bound splits program paths into ranges only entering the loop at most times and ranges entering loops for more than times222Such splits can also be performed on intervals on loop bounds, thereby generating more than two path ranges..
@@
This paper builds on and extends our previous publication [14]. The extension concerns the novel integration of work stealing into ranged analysis and the joining of witnesses. We have furthermore significantly extended our evaluation as to see which improvements the novel concepts bring and as to validate witness joining. Our experiments build on the result of the evaluation described in [14]; we in particular do not repeat the experiments for different splitting strategies as the previous experiments have already shown which splitting strategy is the – on average – best one.
@@ 2 Background
We start by introducing some background concerning the syntax and semantics of programs and proceed by defining path ranges as well as the configurable program analyses implemented in CPAchecker.
@@ 2.1 Program Syntax and Semantics
In this paper, we consider simple, imperative programs with a deterministic control-flow with integer variables only (taken from some overall set of variables )333Our implementation supports C programs.. A program is formally modeled by a control-flow automaton (CFA) , where is a subset of the program locations (the program counter values), is the initial location, and control-flow edges describe the execution of operations in locations leading to next locations. Therein the set of operations comprises assume statements (Boolean expressions over variables , denoted by ), assignments, skip (empty) operations and assertions. We expect that CFAs originate from program code and, thus, control-flow may only branch at assume operations, i.e., CFAs are deterministic in the following sense: For all either or are assume operations and . We assume that there exists an indicator function that reports the branch direction, either N(one), T(rue), or F(alse). This indicator function assigns to all edges without assume operations and for any two assume operations with it guarantees . Since CFAs are typically derived from programs and assume operations correspond to the two evaluations of conditions of e.g., if or while statements, the assume operation representing the true evaluation of the condition is typically assigned . We will later need this indicator function for defining path orderings.
Figure 1 (left) shows our example program, which either iteratively increments two variables or directly assigns the constant value to them. For each condition of an if or while statement the CFA (shown in the middle) contains one assume edge for each evaluation of the condition, depicted as solid edges labeled by the condition for entering the if branch or loop when the condition evaluates to true and dashed edges labeled by the negated condition for entering the else branch or leaving the loop when the condition evaluates to false. All other statements are represented by a single edge. The assertion assert(a==b) at the end of the program is the property to be checked for the program, and here we see that the program is correct.
Next, we give the semantics of programs. A program state is a pair of a program location and a data state from the set of data states that assign to each variable a value of the variable’s domain. Program execution paths are sequences of states and edges such that (1) they start at the beginning of the program and (2) only perform valid execution steps that (a) adhere to the control-flow, i.e., , and (b) properly describe the effect of the operations, i.e., , where the strongest postcondition is a partial function modeling the effect of operation on data states. Execution paths are also called feasible paths, and paths that fulfil properties (1) and (2a) but violate property (2b) are called infeasible paths. The set of all feasible execution paths of a program is denoted by .
@@ 2.2 Path Ordering, Execution Trees, and Ranges
Ranged symbolic execution runs on so called path ranges, which are sets of “consecutive” paths. To specify these sets, we first define an ordering on (feasible) execution paths. Given two program paths and , we define their order based on their control-flow edges. More specifically, edges with assume operations representing a true evaluation of a condition are smaller than the edges representing the corresponding false evaluation of that condition. Following this idea, if . An execution tree is a tree containing all feasible execution paths of a program with the previously defined ordering, where nodes are labeled with the assume operations. We depict a shortened version of the execution tree for our example program in Fig. 1.
Based on the above ordering, we now specify ranges, which describe sets of “consecutive” program execution paths analyzed by a ranged analysis and which are characterized by a left and right path that limit the range. Hence, a range is the set 444In [10], the range is formalized as but their implementation works on .. To easily describe ranges that are not bound on the left or right, we use the special paths which are smaller and greater than every path, i.e., . Consequently, .
As the program is assumed to be deterministic except for the input, a test case , , which maps each input variable to a concrete value, describes exactly a single path 555More concretely, test input describes a single maximal path and all its prefixes.. We say that induces and write this path as . Consequently, we can define a range by two induced paths, i.e., as for two test cases and . For the example program from Fig. 1, two example test cases are and . Two such induced paths are
@@ 2.3 Configurable Program Analysis
To run the configured analysis, one executes a meta reachability analysis, the so-called CPA algorithm, configured by the CPA. To this end, an initial value which the analysis will start with has to be given. For details on the CPA algorithm, we refer the reader to [15].
Value Analysis. As part of our ranged analysis, we use the abstract domain and transfer relation of a CPA for value analysis [16] (also known as constant propagation or explicit analysis). An abstract state of the value analysis ignores program locations and maps each variable to either a concrete value of its domain or , which represents any value. The partial order and the join operator are defined variable-wise while ensuring that 666Consequently, . and if and otherwise . The concretization of abstract state contains all concrete states that agree on the concrete variable values, i.e., . If the values for all relevant variables are known, the transfer relation will behave like the program semantics. Otherwise, it may overapproximate the executability of a CFA edge and may assign value if a concrete value cannot be determined.
Composite CPA. To easily build ranged analysis instances for various program analyses, we modularize our ranged analysis into a range reduction (see Sec. 3.2) and a program analysis. Technically, we will compose a ranged analysis from different CPAs using the concept of a composite CPA [15]. Here, we just define the composition of two CPAs. The composition of more than two CPAs works analogously or can be achieved by recursively composing two (composite) CPAs. A composite CPA of CPA and CPA considers the product domain that defines the operators elementwise, i.e., for pairs of abstract states we define if and , , and . The transfer relation may be the product transfer relation or may strengthen the product transfer relation using knowledge about the other abstract successor. In contrast, and cannot be derived and must always be defined.
@@ 3 Ranged Program Analysis
In this section, we introduce ranged program analysis for generalizing the concept of ranged symbolic execution to arbitrary program analyses. Our goal is to split the set of program paths in a given program into multiple disjoint ranges, each of which can be analysed by an independent program analysis. For the analysis of a given range, we introduce the concept of ranged analysis. A ranged analysis is a program analysis whose exploration is limited to a specific range. We realize ranged analysis as a composition of an arbitrary analysis and a range reduction, both defined as a configurable program analysis (CPA). Then, to analyze the complete program with a ranged program analysis, we (1) split the program into ranges, (2) run several ranged analyses in parallel and (3) aggregate the analysis results. A conceptual overview of the ranged program analysis of two ranges is shown in Fig. 2.
reductions
@@ 3.1 Ranged Analysis
We decompose the range reduction for into a composition of two specialized ranged reductions and , which decide whether a path is in the range and , respectively. Since and the composition stops the exploration of a path if one analysis returns (i.e., no states to explore), the composite analysis only explores paths that are included in both ranges (which are exactly the paths in ). Figure 3 depicts the application of range reduction to the example from Fig. 1, where the range reduction is depicted in Fig. 3a and in Fig. 3b and the composition of both range reductions in Fig. 3c. Finally, to transform any arbitrary program analysis into a ranged analysis, we construct the following composition for a given range :
For , we define component-wise for the individual merge operators and as conjunction of the individual stop operators. As soon as the range reduction decides that a path is not contained in the range and returns , the exploration of the path stops for all analyses defined in the composition.
@@ 3.2 Range Reduction as CPA
Initialisation. To define the CPAs for and , we reuse components of the value analysis (as described in Sec. 2.3). A value analysis explores at least all feasible paths of a program by tracking the values for program variables. If the program behaviour is fully determined (i.e., all (input) variables are set to constants), then only one feasible, maximal path exists, which is explored by the value analysis. We exploit this behaviour by initializing the analysis based on our test case (being a lower or upper bound of a range):
In this case, all variables which are typically undetermined777We assume that randomness is controlled only through an input and hence the program is deterministic. and dependent on the program input have now a determined value, defined through the test case. As the behaviour of the program under the test case is now fully determined, the value analysis only explores a single path , which corresponds to the execution trace of the program given the test case. Now, as we are interested in all paths defined in a range and not only a single path, we adapt the value analysis as follows:
Lower Bound CPA. For the CPA range reduction we borrow all components of the value analysis except for the transfer relation . The transfer relation is defined as follows:
@@
Upper Bound CPA. For the CPA range reduction we again borrow all components of the value analysis except for the transfer relation . The transfer relation is defined as follows:
@@ 3.3 Handling Underspecified Test Cases
So far, we have assumed that test cases are fully specified, i.e., contain values for all input variables, and the behaviour of the program is deterministic such that executing a test case follows a single (maximal) execution path . However, in practice, we observe that test cases can be underspecified such that a test case does not provide concrete values for all input variables. We denote by the set of all paths that are then induced by . In this case, we define:
@@ 3.4 Splitting
To be able to apply ranged program analysis, generating ranges, the splitting of programs into parts that can be analyzed in parallel, is a crucial part. For splitting, a range can either be defined by two program paths or two test cases that need to be computed. We propose and evaluate four different splitting in [14]. The evaluation shows that the strategy called Lb3 generates ranges leading to the best performance of ranged program analysis. Thus, we will only explain Lb3 and refer the reader to [14] for an overview of all strategies.
The Splitter Lb3. The goal of this splitter is to generate a range that contains only three loop iterations. More precisely, the splitter Lb3 computes the left-most path in the program that contains exactly three unrollings of the loop. If the program contains nested loops, each nested loop is unrolled for three times in each iteration of the outer loop. To generate the test case for the computed path, we (1) build its path formula using the strongest post-condition operator [17], (2) use an SMT-solver to check the formula for satisfiability and (3) in case of an answer “SAT”, use the evaluation of the input variables provided by the SMT-solver in the path formula as one test case. In case the path formula is unsatisfiable, we shorten the path by iteratively removing the last statement from it, until we get a satisfying path formula. A test case generated by Lb3 defines the two ranges and . For loop-free programs, we cannot bind the number of loop unrollings. Hence, Lb3 fails and we generate a single range . In our running example in Fig. 1, Lb3 generates .

### v2

@@ Abstract
Symbolic execution is a software verification technique symbolically running programs and thereby checking for bugs. Ranged symbolic execution performs symbolic execution on program parts, so-called path ranges, in parallel. Due to the parallelism, verification is accelerated and hence scales to larger programs.
In this paper, we discuss a generalization of ranged symbolic execution to arbitrary program analyses. More specifically, we present a verification approach that splits programs into path ranges and then runs arbitrary analyses on the ranges in parallel. Our approach in particular allows to run different analyses on different program parts. We have implemented this generalization on top of the tool CPAchecker and evaluated it on programs from the SV-COMP benchmark. Our evaluation shows that verification can benefit from the parallelization of the verification task, but also needs a form of work stealing (between analyses) to become efficient.
@@ 1 Introduction
In recent years, automatic software verification has become a mature technique reaching levels of industrial software. This progress is achieved by verification tools employing a bunch of different techniques for analysis, like predicate abstraction, bounded model checking, k-induction, property-directed reachability, or automata-based methods. As individual tools still have their specific strength and weaknesses, recent research in software verification has studied cooperative verification [1].
In cooperative verification, tools together work on verification tasks, typically by verifying different parts of the software. This principle has already been implemented in various forms [2, 3, 4, 5], in particular also as cooperations of testing and verification [6, 7, 8, 9]. Cooperative verification often takes the form of a sequential (or cyclic) combination in which tools run in some sequential order, passing information from one to the next tool. In such a sequential combination, the first tool might try to verify the entire program, and if it does not succeed, passes information about the unverified part to the next tool, which continues verification. An example of this is conditional model checking [10, 8, 3] in which the first tool passes information about the unverified part in the form of a so-called condition. A condition has an automaton-like structure and can encode (parts of) programs. In such approaches, the division of work (i.e., the split of the program) is not apriori computed, but instead implicitly determined by the first tool.
@@
Parallel combinations (in which different tools operate in parallel) in the majority of cases employ portfolio approaches, simply running different tools on the same task in parallel. In a portfolio approach, there is no need for splitting programs: every tool gets exactly the same program, and typically the first tool completing verification determines the outcome of the verification task. It is even possible to run several instances of the same tool in parallel, just with different configurations. The reason for mainly employing portfolio approaches for parallel combinations is often the lack of appropriate splitting techniques, which could divide the program into parts on which analyses could be run in parallel.
@@
The proposal of Siddiqui and Khurshid [11] for ranged symbolic execution provided exactly one such splitting technique. They defined so-called path ranges as parts of the program on which several symbolic execution instances can be run in parallel. A path range describes a set of program paths defined by two inputs to the program, where the path triggered by the first input is the lower bound and the path for the second input is the upper bound of the range. All paths in between, i.e., paths such that (based on some ordering on paths), make up a range. Then, several symbolic execution engines can be run in parallel on different ranges, searching for bugs in path ranges. Siddiqui and Khurshid experimentally confirmed that symbolic execution can benefit from this splitting into path ranges and this helps to scale the technique to larger programs.
@@
In this paper, we generalize ranged symbolic execution to arbitrary analyses. Our generalization concerns the usage of different analyses. Instead of its specialization to one analysis, namely symbolic execution, we present ranged program analysis in which arbitrary analyses (in particular also different ones) can be run on the different ranges of a program in parallel. In this, we employ a form of work stealing in which one analysis can – once done with its own range – take over ranges assigned to other analyses to better balance workloads. We have furthermore extended the approach of [11] by a new splitting technique. Instead of only choosing ranges randomly, we propose a splitting technique along loop bounds of arbitrary, user-configurable numbers . A loop bound splits program paths into ranges only entering the loop at most times and ranges entering loops for more than times333Such splits can also be performed on intervals on loop bounds, thereby generating more than two path ranges..
@@
This paper builds on and extends our previous publication [15]. The extension concerns the novel integration of work stealing into ranged analysis and the joining of witnesses. We have furthermore significantly extended our evaluation to see which improvements the novel concepts bring and to validate witness joining. Our experiments build on the result of the evaluation described in [15]; we in particular do not repeat the experiments for different splitting strategies as the previous experiments have already shown which splitting strategy is the – on average – best one.
@@ 2 Background
The task in software verification is to decide, whether a given program satisfies a certain property. In this work, we aim to check if a C program can violate an assertion. More precisely, the analyses aim to show the unreachability of an assertion failure. To be able to formally analyze programs, we start by introducing some background concerning the syntax and semantics of programs and proceed by defining path ranges as well as the configurable program analyses implemented in CPAchecker.
@@ 2.1 Program Syntax and Semantics
In this paper, we consider simple, imperative programs with a deterministic control-flow with integer variables only (taken from some overall set of variables )444Our implementation supports C programs.. A program is formally modeled by a control-flow automaton (CFA) , where is a subset of the program locations (the program counter values), is the initial location, and control-flow edges describe the execution of operations in locations leading to next locations. Therein, the set of operations comprises assume statements (Boolean expressions over variables , denoted by ), assignments, skip (empty) operations, and assertions. We expect that CFAs originate from program code and, thus, control-flow may only branch at assume operations, i.e., CFAs are deterministic in the following sense: For all either or are assume operations and . We assume that there exists an indicator function that reports the branch direction, either N(one), T(rue), or F(alse). This indicator function assigns to all edges without assume operations and for any two assume operations with it guarantees . Since CFAs are typically derived from programs and assume operations correspond to the two evaluations of conditions of e.g., if or while statements, the assume operation representing the true evaluation of the condition is typically assigned . We will later need this indicator function for defining path orderings.
Figure 1 (left) shows our example program, which either iteratively increments two variables or directly assigns the constant value to them. For each condition of an if or while statement, the CFA (shown in the middle) contains one assume edge for each evaluation of the condition, depicted as solid edges labeled by the condition for entering the if branch or loop when the condition evaluates to true and dashed edges labeled by the negated condition for entering the else branch or leaving the loop when the condition evaluates to false. All other statements are represented by a single edge. The assertion assert(a==b) at the end of the program is the property to be checked for the program, and here we see that the program is correct.
Next, we give the semantics of programs. A program state is a pair of a program location and a data state from the set of data states that assign to each variable a value of the variable’s domain. Program execution paths are sequences of states and edges such that (1) they start at the beginning of the program and (2) only perform valid execution steps that (a) adhere to the control-flow, i.e., , and (b) properly describe the effect of the operations, i.e., , where the strongest postcondition is a partial function modeling the effect of operation on data states. Execution paths are also called feasible paths, and paths that fulfill properties (1) and (2a) but violate property (2b) are called infeasible paths. The set of all feasible execution paths of a program is denoted by .
@@ 2.2 Path Ordering, Execution Trees, and Ranges
Ranged symbolic execution runs on so-called path ranges, which are sets of “consecutive” paths. To specify these sets, we first define an ordering on (feasible) execution paths. Given two program paths and , we define their order based on their control-flow edges. More specifically, edges with assume operations representing a true evaluation of a condition are smaller than the edges representing the corresponding false evaluation of that condition. Following this idea, if . Note that the path order is a total function, as each feasible execution path of a CFA P starts in .
An execution tree is a tree containing all feasible execution paths of a program with the previously defined ordering, where nodes are labeled with the assume operations. We depict a shortened version of the execution tree for our example program in Fig. 1.
Based on the above ordering, we now specify ranges, which describe sets of “consecutive” program execution paths analyzed by a ranged analysis and which are characterized by a left and right path that limits the range. Hence, a range is the set 555In [11], the range is formalized as but their implementation works on .. To easily describe ranges that are not bound on the left or right, we use the special paths which are smaller and greater than every path, i.e., . Consequently, .
As the program is assumed to be deterministic except for the input, a test case , , which maps each input variable to a concrete value, describes exactly a single path 666More concretely, test input describes a single maximal path and all its prefixes.. We say that induces and write this path as . Consequently, we can define a range by two induced paths, i.e., as for two test cases and . For the example program from Fig. 1, two example test cases are and . Two such induced paths are
@@ 2.3 Configurable Program Analysis
To run the configured analysis, one executes a meta-reachability analysis, the so-called CPA algorithm, configured by the CPA. To this end, an initial value which the analysis will start with has to be given. For details on the CPA algorithm, we refer the reader to [16].
Value Analysis. To restrict our ranged analysis to a particular program range, we track the induced paths defining the left and right border of the range using the abstract domain and transfer relation of a CPA for value analysis [17] (also known as constant propagation or explicit analysis). An abstract state of the value analysis ignores program locations and maps each variable to either a concrete value of its domain or , which represents any value. The partial order and the join operator are defined variable-wise while ensuring that 777Consequently, . and if and otherwise . The concretization of abstract state contains all concrete states that agree on the concrete variable values, i.e., . If the values for all relevant variables are known, the transfer relation will behave like the program semantics. Otherwise, it may overapproximate the executability of a CFA edge and may assign value if a concrete value cannot be determined.
Composite CPA. To easily build ranged analysis instances for various program analyses, we modularize our ranged analysis into a range reduction (see Sec. 3.2) and a program analysis. Technically, we will compose a ranged analysis from different CPAs using the concept of a composite CPA [16]. Here, we just define the composition of two CPAs. The composition of more than two CPAs works analogously or can be achieved by recursively composing two (composite) CPAs. A composite CPA of CPA and CPA considers the product domain that defines the operators elementwise, i.e., for pairs of abstract states we define if and , , and . The transfer relation may be the product transfer relation or may strengthen the product transfer relation using knowledge about the other abstract successor. Similarly, we may define elementwise, i.e., or define a customized, more precise merge operator. While a conjunction of the individual termination checks is typically not sound, in practice most composite analyses use the termination check , which can be automatically derived from the partial order.
@@ 3 Ranged Program Analysis
In this section, we introduce ranged program analysis for generalizing the concept of ranged symbolic execution to arbitrary program analyses. Our goal is to split the set of program paths in a given program into multiple disjoint ranges, each of which can be analyzed by an independent program analysis. For the analysis of a given range, we introduce the concept of ranged analysis. A ranged analysis is a program analysis whose exploration is limited to a specific range. We realize ranged analysis as a composition of an arbitrary analysis and a range reduction, both defined as a configurable program analysis (CPA). Then, to analyze the complete program with a ranged program analysis, we (1) split the program into ranges, (2) run several ranged analyses in parallel, and (3) aggregate the analysis results. A conceptual overview of the ranged program analysis of two ranges is shown in Fig. 2. Next, we formalize ranged analysis in Sec. 3.1. Then, we define in Sec. 3.2 the range reduction, which is used to restrict an arbitrary program analysis to the paths within a range. Thereafter, we explain how to handle underspecified test cases in Sec. 3.3 and introduce the splitter Lb3 for generating ranges in Sec. 3.4.
@@ 3.1 Ranged Analysis
We decompose the range reduction for into a composition of two specialized ranged reductions and , which decide whether a path is in the range and , respectively. Since and the composition stops the exploration of a path if one analysis returns (i.e., no states to explore), the composite analysis only explores paths that are included in both ranges (which are exactly the paths in ).
reductions
Figure 3 depicts the application of range reduction to the example from Fig. 1, where the range reduction is depicted in Fig. 3a and in Fig. 3b and the composition of both range reductions in Fig. 3c. Finally, to transform any arbitrary program analysis into a ranged analysis, we construct the following composition for a given range :
For , we use and define component-wise for the individual merge operators as explained in Sec. 2.3.
@@ 3.2 Range Reduction as CPA
Initialization. To define the CPAs for and , we reuse components of the value analysis (as described in Sec. 2.3). A value analysis explores at least all feasible paths of a program by tracking the values for program variables. If the program behavior is fully determined (i.e., all (input) variables are set to constants), then only one feasible, maximal path exists, which is explored by the value analysis. We exploit this behavior by initializing the analysis based on our test case (being a lower or upper bound of a range):
In this case, all variables which are typically undetermined888We assume that randomness is controlled only through an input and hence the program is deterministic. and dependent on the program input have now a determined value, defined through the test case. As the behavior of the program under the test case is now fully determined, the value analysis only explores a single path , which corresponds to the execution trace of the program given the test case. Now, as we are interested in all paths defined in a range and not only a single path, we adapt the value analysis as follows:
Lower Bound CPA. For the CPA range reduction , we borrow all components of the value analysis except for the transfer relation . The transfer relation is defined as follows:
@@
Upper Bound CPA. For the CPA range reduction , we again borrow all components of the value analysis except for the transfer relation . The transfer relation is defined as follows:
@@ 3.3 Handling Underspecified Test Cases
So far, we have assumed that test cases are fully specified, i.e., contain values for all input variables, and the behavior of the program is deterministic such that executing a test case follows a single (maximal) execution path . However, in practice, we observe that test cases can be underspecified such that a test case does not provide concrete values for all input variables. We denote by the set of all paths that are then induced by . In this case, we define:
@@ 3.4 Splitting
To be able to apply ranged program analysis, generating ranges, the splitting of programs into parts that can be analyzed in parallel, is a crucial part. For splitting, a range can either be defined by two program paths or two test cases that need to be computed. We propose and evaluate four different splitting in [15]. The evaluation shows that the strategy called Lb3 generates ranges leading to the best performance of ranged program analysis. Thus, we will only explain Lb3 in detail and only briefly summarize the other three.
The Splitter Lb3. The goal of this splitter is to generate a range that contains only three loop iterations. More precisely, the splitter Lb3 computes the left-most path in the program that contains exactly three unrollings of the loop. If the program contains nested loops, each nested loop is unrolled three times in each iteration of the outer loop. To generate the test case for the computed path, we (1) build its path formula using the strongest post-condition operator [18], (2) use an SMT-solver to check the formula for satisfiability and (3) in case of an answer “SAT”, use the evaluation of the input variables provided by the SMT-solver in the path formula as one test case. In case the path formula is unsatisfiable, we shorten the path by iteratively removing the last statement from it, until we get a satisfying path formula. A test case generated by Lb3 defines the two ranges and . For loop-free programs, we cannot bind the number of loop unrollings. Hence, Lb3 fails and we generate a single range . In our running example in Fig. 1, Lb3 generates .
Other splitting strategies. Beneath Lb3 we also introduce the splitting strategy Lb10, which is also based on loop unrollings. In contrast to Lb3, Lb10 computes the left-most path that contains exactly ten loop unrollings. The other two strategies select paths randomly by traversing the CFA and deciding at each assume edge to either follow the true- or the false-branch. Rdm selects both branches with a probability of %, whereas Rdm9 selects the true-branch with a % probability. As execution trees of programs with loops are often not balanced but rather grow to the left, Rdm9 likely generates longer paths. For a more detailed description, we refer the reader to [15].